# MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation

### v3

@@ 3.3 ECG Instruction Tuning
where $j$ is the start index after <assistant>, $\theta$ is the trainable parameters of LoRA and ECG encoder $\mathcal{F}_{e}$, $\mathbf{X}_{t,<i}$ is the response tokens before the current generation $\mathbf{x}_{t,i}$.
4 Benchmarking ECG Report Generation
In this section, we create a benchmark to evaluate the ECG report generation approach. Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation, (2) the zero-shot ability under domain transfer, (3) the robustness of the model against perturbed signals.
@@ 4.1 Datasets
PTB-XL. The PTB-XL dataset [11] comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients. Each ECG is accompanied by a corresponding report. The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds. We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively. The human experts double-check all samples in the test set to ensure data quality. As mentioned in Sec 3.1, we reformulate the dataset into the instruction data format.
MIMIC-IV-ECG. The MIMIC-IV-ECG dataset [10] is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects. Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds. The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio. Likewise, we reconstruct this dataset into an ECG instruction data template.
@@ 4.2 Evaluation Metrics
We evaluate models using nine metrics: BLEU 1-4 [17], METEOR [18], ROUGE 1-2 and L [19], CIDEr-D [20], and BERTScore [21]. BLEU and METEOR assess machine translation quality, focusing on accuracy and fluency. ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps. CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
4.3 Tasks
We evaluate the performance of ECG report generation on the following three tasks:
Quality Report Generation. This task aims to assess report quality after ECG instruction tuning using 10% of MIMIC-IV-ECG and PTB-XL datasets as the test set. The evaluation examines how closely generated reports match the original’s structure and meaning, considering various instructions and ECG inputs. We analyze metrics like BLEU-1 to 4, METEOR, ROUGE 1, 2, L, CIDEr-D, and BERTScore.
Zero-shot Generalizability. To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG. Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set. It’s important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents—Europe and the United States, respectively—utilizing varied devices and from distinct hospitals, across different time periods. Therefore, we consider these datasets to represent two separate domains. This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively. We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
Signal Perturbation Robustness. In real-world clinical settings, ECG signals often contain some degree of noise. To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset. We then added Gaussian noise to these samples during the models’ instruction-based inference process. For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
5 Experiments and Evaluation
5.1 Experimental Settings
In this section, we evaluate and benchmark nine open-source decoder-only LLMs using the constructed ECG report generation benchmark. Additionally, we offer a comprehensive analysis of scalability and instruction tuning and present case studies showcasing the generated reports.
Models. We use nine LLMs based on the peft [^1] https://github.com/huggingface/peft library, which directly supports LoRA [16] to construct the multimodal ECG report generation model described in Section 3.2. These models include GPT-Neo [22], GPT-NeoX [23], GPT-J [24], BLOOM [25], OPT [26], LLaMA-1 [27], LLaMA-2 [6], Mistral [28], and Mistral-Instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with
Table 1: Natural language generation metric on MIMIC-IV-ECG. For model size, ’M’ denotes the million level, and ’B’ denotes the billion level. † refers to an improved instruct fine-tuned version of models. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
@@ 5.1 Experimental Settings
LLaMA-2† 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
Mistral-Instruct† 7B 0.714 0.665 0.619 0.576 0.768 0.751 0.762 0.667 5.62
@@
LLaMA-2† 7B 0.515 0.484 0.469 0.439 0.675 0.594 0.698 0.624 4.05
Mistral-Instruct† 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
two relatively small pre-trained language models (GPT2-Medium and GPT-Large [29]) as fundamental baselines.
Implementation Details. Our study utilized PyTorch 2.1, transformers [30], and accelerated on A100
@@
LLaMA-2† 0.764 0.725 0.744 0.721 0.668 0.693
Mistral-Instruct† 0.773 0.722 0.747 0.730 0.661 0.694
GPUs with LLMs from Hugging Face [31] ranging from 2.7 to 70 billion parameters. For larger models, we used DeepSpeed [^3] https://github.com/microsoft/DeepSpeed. The training covered 5 epochs on MIMIC-IV-ECG and PTB-XL with a 2e-5 learning rate and 64 batch size, employing a linear optimizer with a 0.03 warm-up ratio. For text preprocessing, we initially remove all instances of the ‘nan’ string and sentences that consist solely of numerical values. Subsequently, we discard any samples whose reports contain fewer than 5 tokens. For, ECG encoder, we adopt random initialization. Additionally, the default number of generated prompts from GPT-4 is 256, more training, visualization details about ECG instruction tuning are illustrated in Appendix A.2, Section  5.3, and Appendix A.4.
5.2 Benchmark Task Evaluation
5.2.1 Quality Evaluation
Performance on MIMIC-I V-ECG. Table 3 and 1 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG. The results show that all LLMs perform better than smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics. Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics. For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
@@ 5.2.1 Quality Evaluation
The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information. This highlights the significant potential of LLMs in medical signal-to-text generation. Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
Performance on PTB-XL. As shown in Table 2, the models exhibit reduced performance on PTB-XL compared to MIMIC-IV-ECG, which is attributable to the smaller scale of the instruction data in PTB-XL. This underscores the importance of data scale in enhancing instruction-based ECG report generation. Moreover, similar to the MIMIC-IV-ECG results, all LLM-based models show significant improvement over SLMs. Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score. The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
5.2.2 Zero-shot Evaluation in Domain Transfer.
@@
In Figure 3, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset). The assessed models include BLOOM, OPT, LLaMA-1, and Mistral. Firstly, all selected LLMs undergo instruction tuning on the MIMIC-IV-ECG train set, followed by zero-shot testing on the PTB-XL test set verified by human experts, denoted as Zero-shot IT. We also measure the performance of each model in report generation without prior ECG-specific instruction tuning, denoted as Zero-shot W/O IT. PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set. Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the model’s ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT. The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
5.2.3 Robust Analysis with Perturbed ECG Signal.
In a noise stress evaluation [32], we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness. Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG
Figure 4: Signal perturbation robustness analysis on various LLMs
@@
training set and tested on corresponding noise-added signals from its test set. The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR decreased, highlighting the significant interference of ECG noise. Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets. The results demonstrate Mistral’s strong resistance to perturbations. Even with more severe noise, it maintained robustness regarding ROUGE-L and METEOR metrics. Developing an even more robust framework is a goal for future research.
@@
5.3 Analysis
Analysis of ECG Modality Alignment. To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA [33] and additional trainable cross-attention layer in Flamingo [12]. For straightforward input, we follow the design of LLaVA by directly concatenating the ECG encoder’s output embeddings with the sentence’s embeddings before inputting them into the LLM backbones. For the second comparison method, we follow Flamingo by adding a trainable cross-attention layer within the attention block. From Table 4, we observe that the Concatenated-fusion method outperforms the trainable cross-attention method of Flamingo in most metrics and is consistently superior to the Straightforward input method of LLaVA. Consequently, the concatenated fusion is more effective for the LLM backbone’s alignment with fine-grained ECG patterns without necessitating additional trainable parameters.
Figure 5: Instruction tuning loss and METEOR visualization.
@@ 5.3 Analysis
However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant. For example, the F-1 score for the 70B model on the PTB-XL dataset exhibits a marginal increase of 0.02 over the 13B model. Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model. Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance [34].
@@
Ablation Study on ECG Instruction Tuning. We conducted an ablation study to evaluate instruction tuning’s impact on aligning ECG signals with report representations. Utilizing LLMs such as BLOOM, OPT, LLaMA-1, and Mistral without instruction tuning, we allowed direct learning from ECG signals. The findings, illustrated in Figure 7, indicate a significant performance drop across all metrics without instruction tuning, particularly in Mistral. This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning [35].

### v4

@@ 3.5 Instruction Tuning
where $j$ is the start index after <assistant>, $\theta$ is the trainable parameters of LoRA and ECG encoder $\mathcal{F}_{e}$, and $\mathbf{X}_{t,<i}$ is the response tokens before the current generation $\mathbf{x}_{t,i}$.
4 ECG Report Generation Benchmark
@@ 4.1 Datasets
MIMIC-IV-ECG. MIMIC-IV-ECG (Gow et al., ) is the largest publicly available 12-lead ECG dataset. MIMIC-IV-ECG contains $800,035$ samples from $161,352$ subjects. Similar to PTB-XL, each sample includes a raw ECG recording and its report, sampled at $500$Hz for 10 seconds. The dataset is split into training, validation, and test sets at a ratio of 80%:10%:10% respectively.
PTB-XL. We also use the PTB-XL dataset (Wagner et al., 2020), which contains $21,837$ clinical 12-lead ECG recordings collected from $18,885$ patients. Each ECG recording is $10$ seconds long sampled at $500$Hz and has a corresponding report. We divide PTB-XL into training, validation, and test sets at a ratio of 70%:10%:20%, respectively.
4.2 Models
We use two smaller language models and ten LLMs based on the PEFT [^1] https://github.com/huggingface/peft library, which directly supports LoRA (Hu et al., 2021) to construct the multimodal ECG report generation model described in Section 3.4. The two smaller language models are GPT2-Medium and GPT-Large (Radford et al., 2019). The ten LLMs are GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), GPT-J (Wang and Komatsuzaki, 2021), BLOOM (Workshop et al., 2022), OPT (Zhang et al., 2022), LLaMA-1 (Touvron et al., 2023a), LLaMA-2-Instruct (Touvron et al., 2023b), LLaMA-3-Instruct (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and Mistral-Instruct.
@@ 4.3 Evaluation Metrics
We evaluate the performance using ten metrics: BLEU 1-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE 1, 2 and L (Lin, 2004), CIDEr-D (Vedantam et al., 2015), and BERTScore (Zhang et al., 2019). Specifically, BLEU 1-4 and METEOR assess machine translation quality, focusing on accuracy and fluency. ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps. CIDEr-D evaluates the relevance and uniqueness of the generated ECG reports against a candidate set. Lastly, BERTScore assesses semantic similarity to the ground truth.
Table 1: Natural language generation metric on MIMIC-IV-ECG. For model size, ’M’ denotes the million level, and ’B’ denotes the billion level. All checkpoints are downloaded from Hugging Face website. And all models have been fine-tuned using ECG instructions. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
@@
LLaMA-2-Instruct 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
Mistral-Instruct 7B 0.714 0.665 0.619 0.576 0.768 0.751 0.762 0.667 5.62
LLaMA-3-Instruct 8B 0.733 0.686 0.648 0.610 0.799 0.773 0.795 0.686 5.78
@@
LLaMA-2-Instruct 7B 0.515 0.484 0.469 0.439 0.675 0.594 0.698 0.624 4.05
Mistral-Instruct 7B 0.501 0.481 0.457 0.425 0.664 0.592 0.700 0.641 4.01
LLaMA-3-Instruct 8B 0.539 0.513 0.494 0.467 0.698 0.615 0.725 0.646 4.45
4.4 Tasks
Task #1: Quality of the Generated Reports. In our first task, we aim to assess the quality of the generated ECG reports using 10% of PTB-XL and MIMIC-IV-ECG datasets as the test set. Specifically, this task examines how closely the generated reports match the ground truth reports in terms of structure and semantic meaning under various instructions and ECG recordings.
Task #2: Zero-shot Generalizability. As the second task, to explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG. Following this, we evaluated the models’ zero-shot capabilities on the PTB-XL test set. It is important to note that the PTB-XL and MIMIC-IV-ECG datasets originate from different continents, Europe, and the United States, respectively, utilizing different devices and from different hospitals, across different time periods. Therefore, we consider these datasets to represent two separate domains. This distinction allows us to use the PTB-XL dataset to gauge our model’s performance in zero-shot domain transfer effectively. We used the metrics BLEU-4, METEOR, ROUGE-L, and CIDEr-D because of limited space and calculated their average for model evaluation.
Task #3: Signal Perturbation Robustness. In real-world clinical settings, ECG signals often contain some degree of noise. To evaluate the robustness of MEIT against such noisy interference, we selected 10% of the ECG samples from the MIMIC-IV-ECG test dataset. We then added Gaussian noise to these samples during the models’ instruction-based inference process. For this evaluation, we used BLEU-4, METEOR, ROUGE-L, and CIDEr-D as metrics.
Task #4: Evaluation of Alignment with Human Expert Annotations. Lastly, to evaluate the differences between the reports generated by MEIT and human expert annotations, we established specific evaluation criteria and utilized closed-source LLMs to conduct a professional assessment of both the generated reports and expert annotations.
5 Experiments and Analysis
@@
LLaMA-2-Instruct 0.764 0.725 0.744 0.721 0.668 0.693
Mistral-Instruct 0.773 0.722 0.747 0.730 0.661 0.694
LLaMA-3-Instruct 0.798 0.745 0.771 0.745 0.682 0.712
5.1 Experimental Setup
We utilized PyTorch 2.1, transformers (Wolf et al., 2020), and accelerated on A100 GPUs with LLMs from Hugging Face (Wolf et al., 2019) ranging from 2.7 to 70 billion parameters. For larger models, we used DeepSpeed [^2] https://github.com/microsoft/DeepSpeed. The training covered 5 epochs on MIMIC-IV-ECG and PTB-XL with a 2e-5 learning rate and 64 batch size, employing a linear optimizer with a 0.03 warm-up ratio. For text preprocessing, we initially remove all instances of the ‘nan’ string and sentences that consist solely of numerical values. Subsequently, we discard any samples whose reports contain fewer than 5 tokens. For ECG encoder, we adopt random initialization. Additionally, the default number of generated prompts from GPT-4 is 256.
5.2 Quality Evaluation
Performance on MIMIC-IV-ECG. Table 1 and 3 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG. The results show that all ten LLMs perform better than the two smaller language models (GPT2-Medium and GPT2-Large) across all evaluation metrics. Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics. For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore. The observed performance underscores the adeptness of LLMs in generalizing from ECG data showcasing enhanced proficiency in aligning ECG data representations with corresponding textual information. This highlights the significant potential of LLMs in ECG-to-text generation. Particularly, LLaMA-2-Instruct, Mistral-Instruct, and LLaMA-3-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
Performance on PTB-XL. As shown in Table 2, the models exhibit reduced performance on PTB-XL compared to MIMIC-IV-ECG, which is attributable to the smaller scale of the instruction data in PTB-XL. This underscores the importance of data scale in enhancing instruction-based ECG report generation. Moreover, similar to the MIMIC-IV-ECG results, all LLM-based models show significant improvement over SLMs. Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score. The overall experimental results also reveal that Mistral-Instruct, LLaMA-2-Instruct, and LLaMA-3-Instruct are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
Table 4: Evaluation results of LLaMA-2-Instruct and LLaMA-3-Instruct against human expert-annotated ground-truth reports. Each dimension is scored on a scale of 1 to 5.
Model Medical Terminology Accuracy Logical Consistency Completeness Diagnostic Accuracy
LLaMA-2-Instruct 4.25 4.11 3.72 3.60
LLaMA-3-Instruct 4.52 4.38 4.01 3.98
@@ 5.2 Quality Evaluation
5.3 Zero-shot Evaluation in Domain Transfer
@@
Although both PTB-XL and MIMIC-IV-ECG datasets are time-series data, they differ significantly in several aspects, including population (European vs. American), diverse collection devices, continents (Europe vs. US), protocols, and hospitals. These differences introduce substantial medical domain gaps (Bilheimer and Klein, 2010; Ross et al., 2020). In Figure 2, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset). The assessed models include BLOOM, OPT, LLaMA-1, and Mistral. Firstly, all selected LLMs undergo instruction tuning on the MIMIC-IV-ECG train set, followed by zero-shot testing on the PTB-XL test set verified by human experts, denoted as Zero-shot IT. We also measure the performance of each model in report generation without prior ECG-specific instruction tuning, denoted as Zero-shot W/O IT. PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set. Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the model’s ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT. The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
5.4 Robust Analysis with Perturbed ECG Recordings
Figure 3: Signal perturbation robustness analysis on various LLMs.
@@ 5.4 Robust Analysis with Perturbed ECG Recordings
In a noise stress evaluation (Wang et al., 2019), we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness. Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG training set and tested on corresponding noise-added signals from its test set. The results, illustrated in Figure 3, show a performance decline in all LLMs as SNR decreased, highlighting the significant interference of ECG noise. Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets. The results demonstrate Mistral’s strong resistance to perturbations. Even with more severe noise, it maintained robustness regarding ROUGE-L and METEOR metrics.
5.5 Evaluation of Alignment with Human Expert Annotations
In Table 4, we conducted an evaluation of model-generated ECG reports from ECG instruction-tuned versions of LLaMA-2 and LLaMA-3 against 500 ground-truth reports, meticulously annotated by human medical experts. These test annotated data were randomly sampled from the PTB-XL dataset, with all selected reports carefully reviewed and validated by human experts. Each model-generated report was compared with these expert-annotated reports using gpt-4o [^3] https://platform.openai.com/docs/models/gpt-4o, which assessed quality across four dimensions: Medical Terminology Accuracy, Logical Consistency, Completeness, and Diagnostic Accuracy, on a scale of 1 to 5. To evaluate these reports, we employed the following prompt template, which guided GPT-4o’s scoring process across the defined dimensions, as shown in Table 6 in Appendix. This prompt template ensures that GPT-4o evaluates the reports in a structured and consistent manner, highlighting both strengths and weaknesses of the model-generated reports in comparison to human expert annotations. The results indicate that the LLaMA-3 model, with an average Diagnostic Accuracy score of 3.85, closely matches the quality of the human expert annotations, whereas the LLaMA-2 model scored 3.60. This evaluation underscores the effectiveness of using human expert annotations from the PTB-XL (Wagner et al., 2020) dataset as a rigorous benchmark for assessing the models’ ability to generate clinically reliable ECG reports.
6 Conclusion
In this paper, we introduced MEIT, a framework for generating instruction-following data to train a multimodal LLM that can produce ECG reports based on human instructions. We also proposed an effective method for aligning ECG and report representations across various open-source LLMs, demonstrating strong performance on both the MIMIC-IV-ECG and PTB-XL datasets across multiple tasks. Additionally, we established a comprehensive benchmark for ECG instruction-following in report generation, providing a standardized evaluation for future research. Although this work primarily focuses on ECG signals, it serves as a foundational step in applying instruction-tuning to biomedical signals. For future research, we aim to extend our framework and benchmark to other medical domains, such as EEG, with the hope of driving further progress in developing more capable medical-signal LLMs.
7 Limitations
Our MEIT framework first attempts to address automatic ECG report generation through ECG instruction tuning, establishing the first comprehensive benchmark for this process with mainstream LLMs. In this work, we mainly focus on generating reports from multimodal ECG instructions using LLMs. However, the generated results are not fully explainable or controllable, even though the generation procedure is transparent and trackable. This is because the underlying theory of LLMs remains largely unexplored, necessitating further investigation to ensure the quality and safety of the generated content. In the future, we aim to enable LLMs to utilize external, expert-verified knowledge databases, such as clinical protocols and medical textbooks, to enhance the explainability of the generated ECG reports.
8 Acknowledgment
We thank the reviewers and ACs for their helpful comments.
Appendix A Appendix.
A.1 Further Analysis of MEIT
Figure 4: Visualizations of instruction tuning loss and METEOR score.
@@ A.1 Further Analysis of MEIT
Analysis of ECG Modality Alignment. To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA (Liu et al., 2023c) and additional trainable cross-attention layer in Flamingo (Alayrac et al., 2022). For straightforward input, we follow the design of LLaVA by directly concatenating the ECG encoder’s output embeddings with the sentence’s embeddings before inputting them into the LLM backbones. For the second comparison method, we follow Flamingo by adding a trainable cross-attention layer within the attention block. From Table 5, we observe that the Concatenated-fusion method outperforms the trainable cross-attention method of Flamingo in most metrics and is consistently superior to the Straightforward input method of LLaVA. Consequently, the concatenated fusion is more effective for the LLM backbone’s alignment with fine-grained ECG patterns without necessitating additional trainable parameters.
@@
However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant. For example, the F-1 score for the 70B model on the PTB-XL dataset exhibits a marginal increase of 0.02 over the 13B model. Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model. Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance (Wei et al., 2022).
@@
Ablation Study on ECG Instruction Tuning. We conducted an ablation study to evaluate instruction tuning’s impact on aligning ECG signals with report representations. Utilizing LLMs such as BLOOM, OPT, LLaMA-1, and Mistral without instruction tuning, we allowed direct learning from ECG signals. The findings, illustrated in Figure 6, indicate a significant performance drop across all metrics without instruction tuning, particularly in Mistral. This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning (Ouyang et al., 2022).
Table 6: Prompt template used for GPT-4o evaluation. This prompt guided the model’s evaluation of generated ECG reports.
Prompt Template for GPT-4o Evaluation
You are an expert in Electrocardiogram (ECG) text evaluation. Your task is to assess the quality of a generated ECG report by comparing it to a real, expert-annotated ECG report.
Generated ECG Report: {Generated_Report}
Real ECG Report: {Real_Report}
Please evaluate the generated report based on the following criteria:
1. Medical Terminology Accuracy: Does the generated report use correct and appropriate ECG signal terms?
2. Logical Consistency: Is the information presented in a logical and medically sound order?
3. Completeness: Does the report include all necessary details that would be present in a real ECG report, such as heart rhythm, rate, and any abnormalities?
4. Diagnostic Accuracy: Are the diagnoses and interpretations in the generated report accurate and consistent with the expert-annotated report?
Please provide a detailed analysis and score each criterion on a scale of 1 to 5 (1 = Poor, 5 = Expert-Level).